### v2

# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

Abstract
Long-range dependencies are critical for effective graph representation learning, yet most existing datasets focus on small graphs tailored to inductive tasks, offering limited insight into long-range interactions. Current evaluations primarily compare models employing global attention (e.g., graph transformers) with those using local neighborhood aggregation (e.g., message-passing neural networks) without a direct measurement of long-range dependency. In this work, we introduce City-Networks, a novel large-scale transductive learning dataset derived from real-world city road networks. This dataset features graphs with over $10^{5}$ nodes and significantly larger diameters than those in existing benchmarks, naturally embodying long-range information. We annotate the graphs based on local node eccentricities, ensuring that the classification task inherently requires information from distant nodes. Furthermore, we propose a model-agnostic measurement based on the Jacobians of neighbors from distant hops, offering a principled quantification of long-range dependencies. Finally, we provide theoretical justifications for both our dataset design and the proposed measurement—particularly by focusing on over-smoothing and influence score dilution—which establishes a robust foundation for further exploration of long-range interactions in graph neural networks. Our dataset and measurement are now both available on PyTorch Geometric, and the code for reproducing the experimental results can be found at github.com/LeonResearch/City-Networks.
Figure 1: Visualizations of City-Networks for Paris, Shanghai, Los Angeles, and London.
1 Introduction
Graphs are a widely used mathematical abstraction across nearly every branch of science. They are particularly effective for modeling the intricate and non-uniform interactions found in real-world data, where nodes stand in for objects and edges depict their connections. The growing recognition of the versatility of graph representations has sparked intense interest in Graph Neural Networks (GNNs) (Scarselli et al., 2009; Wu et al., 2021), driving innovation in deep learning for both geometric and graph-centric applications. Most GNNs, in particular variants based on a message-passing mechanism (Gilmer et al., 2017), exchange information between one-hop neighbors per layer to build node representations. While they found wide success in analysing citation and social networks (Yang et al., 2016b), one significant limitation concerns their capability of handling long-range dependencies, where interactions between distant nodes might be required to solve a task. Most existing datasets are not sufficient in assessing this: for instance, social networks, despite comprising thousands of nodes, often exhibit the small world effect (Watts & Strogatz, 1998) with short average path lengths and high clustering coefficients. The node labels on these graphs usually possess a homophilic pattern where nodes tend to connect with “similar” or “alike” others (McPherson et al., 2001), making it possible to propagate sufficient information for modelling with only a few message-passing layers. On the other hand, while connected nodes tend to have different properties on heterophilic graphs (Zhu et al., 2020; Ma et al., 2022), solving the tasks in those cases do not necessarily require the handling of long-range dependencies (Arnaiz-Rodriguez & Errica, 2025).
Recently, Long Range Graph Benchmark (LRGB(Dwivedi et al., 2023) introduces alternative graph datasets based on super-pixels and molecules. To justify the existence of long-rangeness, the authors compare classical GNNs with Graph Transformers (GTs), which leverage global attentions over the entire graph, and associate the observed performance gaps with the presence of long-range dependencies. However, conclusions that are solely derived from empirical comparisons may not be reliable as they can be largely influenced by hyperparameter tuning (Tönshoff et al., 2024), leading to an ambiguous assessment of the long-range interactions. Moreover, LRGB and other synthetic long-range benchmarks (Bodnar et al., 2021; Zhou et al., 2025) all focus on inductive learning tasks with relatively small graphs—typically on the order of $10$ to $10^{2}$ nodes, while currently there is no long-range benchmark that considers large graphs with real-world topology for transductive learning. This is a critical gap in the literature since applying GTs (Rampášek et al., 2022), which are expected to better capture long-range interactions, to large graphs is significantly more challenging compared to small-graph inductive tasks due to the computational complexity of global attention (Borde, 2024).
We aim to address these limitations in this work, and our main contributions are as follows:
  • We propose City-Networks, a transductive learning dataset that consists of four large-scale city road networks with a topology distinct from those commonly found in the literature. In particular, it features grid-like large graphs with up to $500k$ nodes and diameters of up to $400$, where the labels are annotated based on node accessibility that inherently requires long-range dependency in its calculation. To the best of our knowledge, this is the first large-graph dataset designed for testing long-range dependencies in graph representation learning.
  • We empirically test classical GNNs and GTs on our dataset under different model depths, and compare their behaviors to those on other common graph datasets that are short-ranged, large-scale, heterophilic, or long-range dependent. The results on our datasets, unlike those on the existing datasets, suggest that communication with neighbors from distant hops consistently improves the performance of all models, supporting the presence of long-range signals.
  • To quantify such long-range dependency, we further introduce a model-agnostic measurement that quantifies per-hop influence of a focal node’s neighbors on its prediction, which is estimated by the aggregated $\ell_{1}$-norm of the Jacobian from a trained model at each hop around the focal node. As a validation, we observe that distant hops exert a greater influence on all baseline models in our City-Networks compared to those on other commonly used graph datasets in the literature.
  • Finally, we theoretically justify the graph structure in our dataset from a spectral perspective on over-smoothing, whose rate we link to the algebraic connectivity and diameter of the graph. In addition, we relate our proposed measurement to the concept of influence as defined in the literature, and study the dilution of mean influence score in grid-like graphs to justify our design.
2 The City-Networks Dataset
In this section, we begin by identifying the limitations and challenges in current literature. Then, we characterize how the topology of City-Networks differs from the existing datasets, and proceed to justify the rationale behind our long-range labeling strategy. Lastly, we explain how our dataset addresses the current limitation and discuss the new challenges brought to the field.
Challenges in testing long-range dependency. To design and fairly evaluate graph datasets of long-range dependency, we need to address three research challenges:
  1. 1.
    How can we generate long-range signals in a principled and controllable manner, so that models are required to communicate with sufficiently distant neighbors of a node to predict its label?
  2. 2.
    Beyond the small graphs used in LRGB or other synthetic benchmarks, how can we design long-range signals on large real-world networks to test the scalability of a model?
  3. 3.
    How can we define a principled measure to quantify and compare the level of long-rangeness across different datasets?
We will discuss how our proposed dataset handles the first two challenges in this section, and address the third challenge in Section 4.
Table 1: Statistics of City-Networks compared to common graph datasets, where $d$, $C$, $T$, diamdiam, homohomo represent degree, clustering coefficient, transitivity, diameter, and homophily, respectively.
Dataset #Nodes #Edges avg($d$) std($d$) max($d$) $C$ $T$ diamdiam homohomo
Paris $114k$ $183k$ $3.2$ $0.8$ $15$ $0.03$ $0.03$ $121$ $0.70$
Shanghai $184k$ $263k$ $2.9$ $1.0$ 88 $0.04$ $0.04$ $123$ $0.75$
Los Angeles $241k$ $343k$ $2.8$ $1.0$ 99 $0.04$ $0.05$ $171$ $0.75$
London $569k$ $759k$ $2.7$ $1.0$ $10$ $0.04$ $0.05$ $404$ $0.76$
Cora $2.7k$ $5.3k$ $3.9$ $5.2$ $168$ $0.24$ $0.09$ 1919 $0.81$
ogbn-arxiv $169k$ $1.16m$ $13.7$ $68.6$ $13k$ $0.23$ $0.02$ $25$ $0.65$
Amazon-Ratings $24k$ $93k$ $7.6$ $6.0$ $132$ $0.58$ $0.31$ $46$ $0.38$
PascalVOC-SP $479$ $1.3k$ $5.7$ $1.2$ $10$ $0.43$ $0.40$ $28$ $0.92$
2.1 Large-scale Road Networks with Long Diameters
Real-world network topologies and features. As shown in Figure 1, our City-Networks consists of street maps in four cities: Paris, Shanghai, Los Angeles, and London, which are obtained by querying OpenStreetMap (Haklay & Weber, 2008) with OSMnx (Boeing, 2024). In particular, we consider a city network inclusive of all types of roads in the city (e.g., drive, bike, walk, etc.), where nodes represent road junctions with features like longitude, latitude, land use, etc.; and edges represent road segments with features like road length, speed limit, road types, etc. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features, where we also make the graphs undirected so that they only represent connectivity between junctions, and retain the largest connected component in each city. We refer readers to Appendix B.1 for a more detailed discussion on our dataset processing approach.
Large graphs with long diameters and low maximum degrees. We can observe from Table 1 that our datasets feature large diameters from $100$ to $400$, which are much longer than those of the common graphs used in the literature. In particular, compared to the super-pixel graphs in PascalVOC-SP with typically $500$ nodes, our datasets have much larger sizes ranging from $100k$ to $500k$. Meanwhile, compared to social networks such as Cora, ogbn-arxiv, and Amazon-Ratings, our City-Networks have smaller average clustering coefficients and much lower maximum degrees. We argue that this distinct graph topology enables us to effectively design learnable long-range signals on graphs, as explained in the next paragraph and further justified theoretically in Section 5.
2.2 Long-range labels based on urban accessibility
Based on the city road networks, we design a real-world task that requires the handling of long-range dependencies. Specifically, we aim to predict how accessible a road junction is based on its own location as well as neighboring characteristics of the urban landscape. This is a useful task especially when it comes to evaluating urban design principles such as the 15-minute city (Abbiasov et al., 2024). To quantify accessibility, we measure the distance one needs to travel from one road junction to its $k$-hop neighbors along road segments in the road network. By design, solving this task requires the model to be able to access information within the $k$-hop neighborhood of each focal node and, by adjusting $k$, we can design a long-range task as desired. Such a notion of accessibility is directly related to node eccentricity $\varepsilon(v)$ in network science (Newman, 2018), which looks at the maximum distance from $v$ to all other nodes in the graph $\mathcal{G}=(V,E)$.
However, using the exact eccentricity as signals is not ideal for two reasons. First, we cannot control the level of long-rangeness, as eccentricity is determined by the entire graph structure and not adjustable by design. In particular, although predicting $\varepsilon(v)$ appears to require distant information based on its definition, in many cases (Figure 6, Appendix B.2), the underlying signal can be well inferred by spatial features (e.g., geographical coordinates) alone. This is because nodes in the city center generally have lower eccentricity due to the grid-like topology in road networks. Second, the calculation of eccentricity requires computing all pair-wise shortest paths in the network, which is typically at cost $\mathcal{O}(|V|^{3})$ and is hence prohibitively expensive for labeling our large-scale networks.
Node classification task based on local eccentricity. With the above consideration in mind, we propose a local eccentricity measure $\hat{\varepsilon}_{k}(v)$ that only considers $k$-hop neighbors $\mathcal{N}_{k}(v)$ of node $v$ when calculating the longest shortest paths with the following expressions:
$\hat{\varepsilon}_{k}(v)=\max_{u\in\mathcal{N}_{k}(v)}~\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}~\sum_{e\in\pi}w(e),\vskip-2.84544pt$ (1)
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ is the set of all possible paths from node $v$ to node $u$, and $w(e)$ denotes the edge weight for edge $e$ (we use road length for this particular purpose, while in all other cases, the graphs are considered unweighted). To design a classification task, we then split $\hat{\varepsilon}_{k}(v)$ for all nodes into $10$ quantiles which we use as node labels for the classification task.
Figure 2: Visualizations of node accessibility estimations based on local eccentricity in two sub-regions, where darker colors indicate smaller eccentricity values, i.e., nodes that are easier to access.
We emphasize that this local approach is introduced not only for computational efficiency, but it also brings in two important considerations. First, it creates long-range signals that extend to $k$-hop neighbors by design, i.e., a known “ground-truth” range, hence is directly controllable and testable. Second, it ensures both graph topology and node features contribute to the modelling, avoiding the problem of high correlation between node signal with either the graph structure or node features alone. Indeed, since $k$ and the edge weights used to compute $\hat{\varepsilon}_{k}(v)$ are unknown to the model, it must integrate structural information from distant neighborhoods with their spatial information (geographic location, road type, land use, etc.) to infer the long-range signal, which makes the task non-trivial.
Choice of the long-range level $k$. To assess long-range dependency, $k$ should be sufficiently large to distinguish our setting from short-range tasks (e.g., social graphs with typical message passing around $4$ hops). Thanks to the large grid-like topology with long diameters, we can create long enough node signals based on different local networks. At the same time, $k$ should also not be too large such that (i) the selected neighborhoods maintain local characteristics; and (ii) baseline models under a $k$-layer architecture, which captures the required information, can fit into common GPUs for fair benchmarking purposes. After experimenting with values from $8$ to $32$ (Figure 8, Appendix B.4), we design the task at kk=$16$, which strikes a balance between sufficient long-rangeness and computational efficiency. We refer readers to Appendix B.4 for a more detailed justification of our methods.
Visualizations and interpretations. As shown in Equation (1), calculating $\hat{\varepsilon}_{k}(v)$ naturally requires information from distant neighbors up to $k$ hops, and it carries a practical meaning which relates to the accessibility of node $v$ in the network. As an illustration, we visualize two sub-regions from the city maps in Figure 2 with kk=$16$. We can observe that nodes on major transportation routes such as freeways and highways tend to have larger eccentricities than those in populated areas. This is because reaching a node’s 16-hop neighbors from a highway junction often requires traveling a much longer distance compared to road junctions in a downtown area. It is also clear from Figure 2 that a significant part of the graph topology is grid-like and possibly quasi-isometric to a lattice, as we will later discuss in Section 6. Note that the homophily scores, as reported in Table 1, are reasonably high since nodes with similar $\hat{\varepsilon}_{k}$ tend to cluster together.
New challenges for learning long-range dependencies. The proposed dataset presents a significant challenge for both GNNs and GTs. In general, such long-range dependencies create a fundamental trade-off for graph learning: while increasing the number of layers in GNNs helps propagate information from distant hops, it also leads to issues such as over-smoothing (Li et al., 2018; Nt & Maehara, 2019; Rusch et al., 2023), over-squashing (Alon & Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and vanishing gradients (Arroyo et al., 2025). On the other hand, GTs, which rely on attention mechanisms with quadratic computational complexity, face scalability challenges when applied to our large-scale city networks compared to learning on smaller long-ranged graphs like those in LRGB (Dwivedi et al., 2023) and other synthetic benchmarks (Zhou et al., 2025).
3 Benchmarks: From Classical GNNs to Graph Transformers
In this section, we benchmark classical GNNs and GTs at different numbers of message-passing layers on City-Networks, and then contrast their behaviors with results on other commonly used graphs to examine the long-range dependencies across datasets.
Experimental setups. We consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all City-Networks, in which we evaluate various classical GNNs and GTs. Specifically, we benchmark GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) for GTs. In addition, an MLP is also included to reflect the importance of graph topology (or lack thereof) in our task. Lastly, we fix the data split for benchmarking purposes and report the mean and standard deviation of the test accuracy over 5 runs with different random seeds.
Figure 3: Baseline results across datasets at different number of layers $L=[2,4,8,16]$. The results for GraphGPS are not shown on London as it is Out-of-Memory on our 4848GB GPU; the result for SGFormer on PascalVOC-SP is also not reported as it’s not originally designed for inductive setting.
Evaluation protocols. As a hypothesis, when a graph model is predicting on a certain node $v$, it will benefit from communicating with $v$’s neighbors from distant hops if the task is long-ranged. Following this rationale, we evaluate baseline models at different numbers of layers $L=[2,4,8,16]$, with the expectation that a larger $L$ would lead to better performance after sufficient training. By comparison, we also evaluate the same baselines on four representative datasets from the literature: Cora, ogbn-arxiv, Amazon-Ratings, and PascalVOC-SP from LRGB, which are homophilous, large-scale, heterophilous, and long-range dependent, respectively. We then compare the behaviors of baseline models across different datasets while increasing the number of layers $L$.
For fair comparison purposes, we follow the latest GNN tuning technique from Luo et al. (2024) that considers residual connection, normalization, dropout, etc. Readers are also referred to Appendix C.2 for full details on our training and experimental setups.
Results. From Figure 3, we can observe a clear and consistent improvement in performance for all baselines on our City-Networks as their message-passing depth $L$ increases from $2$ to $16$. It suggests that, even with sufficient training, a shallow graph model cannot outperform its deeper counterpart if it does not communicate with sufficiently distant hops. This result also means that the gains from incorporating long-range information as $L$ increases outweigh other side effects of depth. By contrast, the performance of baseline models gradually degrades as $L$ increases on Cora and ogbn-arxiv due to the locality of these tasks and depth-related issues. To elaborate on this matter, we further provide a theoretical justification for our dataset in Section 5, where we link the rate of over-smoothing (a prominent issue associated with depth) to the algebraic connectivity and network diameter of the underlying graph. On the other hand, the baseline performance remains relatively unchanged with only a slight increase on Amazon-Ratings and PascalVOC-SP, which suggests that these tasks are relatively short-ranged compared to City-Networks. To examine this hypothesis, we therefore introduce a quantitative measure of long-range dependency in the next section and compare it across datasets and baselines.
For completeness, we further conducted an ablation study in Appendix C.3 that (i) tests baselines under different hidden_size; and (ii) fixes the model size at $L=16$, refrains the model from seeing beyond hop-$H$ at each node, then tests baselines at different $H$. The results are all consistent with our above findings. In addition, we also summarize the best baseline performance on City-Networks in Appendix C.4, and discuss the potential limitation of our experiments in Appendix E.
4 A Long-range Measurement Based on Jacobians
In this section, we propose to quantify the long-range dependency of a task given a trained model, by evaluating how the influence score $I(v,u)$ between node-pairs  (Xu et al., 2018; Gasteiger et al., 2022) varies with distance. The influence score measures the sensitivity of a GNN layer at node $v$ to the input feature of node $u$ using the Jacobian: $I(v,u)=\sum_{i}\sum_{j}\left|\,\partial\bm{H}^{(L)}_{vi}\big/\partial\bm{X}_{uj}\,\right|$, where $\bm{H}^{(L)}_{vi}$ is the $i^{th}$ entry of node $v$’s embedding at layer $L$, and $\bm{X}_{uj}$ is the $j^{th}$ entry of the input feature vector for node $u$. Unless otherwise specified, we assume $L$ refers to the last logit layer.
Based on the influence score, we define the average total influence $\bar{T}_{h}$ at $h^{th}$ hop as:
$T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u),\quad\bar{T}_{h}=\frac{1}{N}\sum_{v}I_{\text{sum}}(v,h),$ (2)
where $T_{h}(v)=I_{\text{sum}}(v,h)$ is the total influence from the $h^{th}$-hop neighbors of node $v$, $\rho(v,u)$ is the length of the shortest path between $v$ and $u$ (note that this is equivalent to the $h$-hop shell later discussed in Section 6), and $N$ is the number of nodes in the network. Also, we would like to highlight that when $h=0$, $T_{0}(v)=I(v,v)$ becomes the influence of the feature at node $v$ on its output. The average total influence quantifies how much, on average, the features of nodes that are $h$ hops away affect the output at the focal node. In other words, by summing the Jacobian‐based influence scores from all $h$-hop neighbors and then averaging over all nodes, it provides an expected measure of the cumulative effect that distant nodes have on each node as the focal node.
Given Equation (2), we further define the average size $R$ of the influence-weighted receptive field as:
$R=\frac{1}{N}\sum_{v\in V}\frac{\sum_{h\geq 0}^{H}h\cdot T_{h}(v)}{\sum_{h\geq 0}^{H}T_{h}(v)},$ (3)
where $H$ is the maximum number of hops to be considered. Intuitively, one can understand $R$ as measuring how far away the average unit of influence is, where graphs with long-range dependencies are expected to have higher proportional influence between more distant nodes. Note that while the proposed $R$ bears some similarity to the recent work (Bamberger et al., 2025), they do not consider per-hop influence $T_{h}(v)$ as a diagnostic tool of the dependency decay. Finally, we refer readers to Appendix D for a discussion on the computational cost of this measurement.
Table 2: Average size of the influence-weighted receptive field $R$ across different datasets and models.
Model Paris Shanghai L.A. London Cora ogbn-arxiv Amazon PascalVOC
GCN 4.86 5.55 5.36 6.09 2.56 1.34 1.92 3.38
GraphSAGE 4.92 5.73 5.44 5.97 2.37 1.40 1.80 2.99
GCNII 3.62 3.64 3.68 3.66 2.76 3.5 1.79 2.60
GraphGPS 8.18 7.88 7.92 OOM 2.65 OOM 6.86 7.14
Exphormer 5.71 7.06 7.15 7.90 2.84 2.80 1.43 1.42
SGFormer 3.75 4.25 4.03 4.01 3.21 1.21 2.46 NA
Results. We validate $R$ and $\bar{T}_{h}$ on City-Networks using the baselines at $L=16$ layers from Section 3. Since PascalVOC-SP is under an inductive setting, we randomly sample 100 graphs from its testing set (more than $400k$ nodes in total) and report their average. The results in Table 2 show that $R$ is consistently higher on our City-Networks across all models compared to those on the other graph datasets, indicating a longer effective range of influence.
Figure 4: Normalized average total influence $\bar{T}_{h}/\bar{T}_{0}$ averaged across nodes at different distances. Note that the influence calculation for GraphGPS is Out-of-Memory on London and ogbn-arxiv.
To better compare $\bar{T}_{h}$ across models and datasets, we normalize it by $\bar{T}_{0}$ and report $\bar{T}_{h}/\bar{T}_{0}$ in Figure 4. For all models, we generally observe a rapid influence drop at distant hops on Cora and ogbn-arxiv, whereas the decay is noticeably much slower on City-Networks. The patterns on Amazon-Ratings and PascalVOC-SP generally fall between the above two cases with influence concentrated on the first few hops, which further corroborate our findings in Section 3. While $R$ and $\bar{T}_{h}$ are both model-dependent, the above results yield a consistent ordering across datasets: models trained on City-Networks are required to leverage information from more distant hops to perform well, compared to those trained on existing datasets. For the rest of the paper, we will first provide a theoretical justification of the proposed dataset by linking over-smoothing to algebraic connectivity and network diameter in Section 5, and then present the intuition behind the proposed long-range measurement in Section 6.
5 Theoretical Justification of Topologies in City-Networks
In this section, we provide a spectral analysis of over-smoothing in GNNs to justify our dataset design. As such, our main goal is not an in-depth theoretical analysis, since both over-smoothing and spectral analysis of GNNs have been carried out extensively in the literature (Rusch et al., 2023; Wu et al., 2023; Oono & Suzuki, 2020; Rong et al., 2019); instead, our aim is to build on top of these and derive simple results which provide a theoretical grounding of our dataset construction. The section is organized as follows. First, we review the concept of information loss in the limit of infinite GNN layers due to over-smoothing, which effectively vanishes the original node features and makes representations collapse to a value that is only dependent on the graph topology (Section 5.1 and Appendix F.1). Next, we relate the rate at which over-smoothing happens and the eigenvalues of the normalized adjacency operator in linearized GCNs (Section 5.2). Given that eigenvalues with larger absolute values slow down over-smoothing, we analyze how the magnitude of the eigenvalues relates to the graph topology of our datasets. This provides justification that sparse graphs with large diameters promote bigger positive eigenvalues for the normalized adjacency operator and hence are less vulnerable to over-smoothing (Theorem 5 and Appendix F.3). Finally, we argue that less over-smoothing makes it possible for GNNs to capture long-range dependencies in the case of our proposed dataset. All proofs are presented in Appendix F.2.
5.1 Preliminaries: A Spectral Perspective on Over-smoothing for Linear GCNs
We begin by analyzing over-smoothing through the lens of linearization, following the framework established in Wu et al. (2019) to build Simple Graph Convolutional Networks (SGCs). We select this model because it is a linearized and more interpretable version of the widely adopted GCN model, which has also been used for theoretical analysis in previous works such as (Giovanni et al., 2022).
Let $\mathcal{G}=(V,E)$ be an undirected graph with $N=|V|$ nodes, adjacency matrix $\bm{A}\in\mathbb{R}^{N\times N}$ and degree matrix $\bm{D}\in\mathbb{R}^{N\times N}$ ($\bm{D}_{ii}=\sum_{j}\bm{A}_{ij}$). The normalized (augmented) adjacency operator $\bm{\tilde{S}}_{adj}\in\mathbb{R}^{N\times N}$ is defined as $\bm{\tilde{S}}_{adj}=(\bm{I}+\bm{D})^{-\frac{1}{2}}(\bm{I}+\bm{A})(\bm{I}+\bm{D})^{-\frac{1}{2}}$, where the original adjacency matrix and degree matrix have been augmented with self-loops. Wu et al. (2019) discuss the spectral properties of the normalized adjacency operator, which satisfies: $\lambda_{N}=1$, and $|\lambda_{i}|<1$ for all $i<N$. Therefore $\lim_{l\to\infty}(\bm{\tilde{S}}_{adj})^{l}=\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}$, where $\bm{\tilde{u}}_{N}$ is the eigenvector corresponding to $\lambda_{N}$, whose entries at node $v$ are $\sqrt{1+\text{degree}(v)}$.
Via linearization (detailed in Appendix F.1), as the number of layers tends to infinity, the graph convolution operation collapses all node features to scalar multiples of $\bm{\tilde{u}}_{N}$, resulting in complete loss of the original feature information, that is, the learned representations suffer from over-smoothing.
5.2 Over-smoothing Rate on Algebraic Connectivity, Diameter, and Sparsity
We now provide intuition about the types of graph topologies that would mitigate the over-smoothing problem. In particular, sparse graphs—those lacking the small-world effect commonly found in citation networks—and graphs with large diameters tend to experience over-smoothing at a slower rate. We argue that this slower rate of over-smoothing implies a higher likelihood that GNNs can learn long-range dependencies during optimization. We validate this intuition by bounding the eigenvalues of the normalized adjacency matrix using graph properties, then showing how these eigenvalues influence over-smoothing, and thus how controlling these graph properties can reduce the likelihood of over-smoothing. More concretely, we extend (Wu et al., 2019, Theorem 1), which only covers $\lambda_{1}$ and $\lambda_{N}$, to Proposition 4 which we prove in Appendix F.2, covering the normalized algebraic connectivity $\lambda_{N-1}$ and showing that $\lambda_{N-1}\left(\bm{S}_{adj}\right)<\lambda_{N-1}(\tilde{\bm{S}}_{adj})$.
Proposition 5.1 (Self-loops decrease algebraic connectivity of the original graph).
Assume a connected graph $\mathcal{G}$ with more than two nodes. For all $\gamma>0$,
$\lambda_{N-1}\left(\bm{S}_{adj}\right)=\lambda_{N-1}\left(\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}\right)<\lambda_{N-1}\left((\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}(\gamma\bm{I}+\bm{A})(\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}\right).$ (4)
This allows us to relate the second largest positive eigenvalue of $\bm{\tilde{S}}_{adj}$ to the topology of graphs in the proposed dataset, using bounds similar to those in (Chung, 1997, Lemma 1.14).
Theorem 5.2 (Bound on second largest positive eigenvalue of the normalized adjacency operator).
Let $d_{max}$ be the maximum degree of a vertex in $\mathcal{G}$, and $diam(\mathcal{G})$ the diameter of $\mathcal{G}$, which must be $diam(\mathcal{G})\geq 4$. Then
$\lambda_{N-1}(\bm{\tilde{S}}_{adj})>\frac{2\sqrt{d_{max}-1}}{d_{max}}-\frac{2}{diam(\mathcal{G})}\left(1+\frac{2\sqrt{d_{max}-1}}{d_{max}}\right).$ (5)
Table 3: The lower bound for $\lambda_{N-1}(\bm{\tilde{S}}_{adj})$ on different datasets.
Dataset Paris Shanghai L.A. London Cora arxiv Ratings Pascal
Lower Bound 0.4741 0.6344 0.6095 0.5921 0.0324 -0.0639 0.1224 0.4857
From Table 3, we see that the lower bound in Equation (5) is generally higher for our datasets; this is because the bound is decreasing in $d_{max}$ and increasing in $diam(\mathcal{G})$, showing graphs with large diameter and low maximum degree will be more resilient to over-smoothing. We further provide the rationale behind this lower bound in Appendix F.3.
6 Theoretical Justification of Jacobian-Based Measurement
Finally, we justify the proposed measurement for analyzing long-range interactions in GNNs, particularly by comparing the mean and the sum of influence scores of neighboring nodes as candidates to represent the total influence score for a focal node $v$, $T_{h}(v)$. Recall that $T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u)$ as in Equation (2). We first introduce definitions for the standard lattice in Euclidean space, $h$-hop shells, and quasi-isometric graphs, and then provide a mathematical analysis of influence score dilution. Due to space limits, we provide the definitions in Appendix H.
Lemma 6.1 (Growth of $h$-hop shells in grid-like graphs).
Let $\mathcal{G}=(V,E)$ be a graph that is grid-like in $D$ dimensions (e.g., let us presume the graphs in the proposed dataset are a subgraph of $\mathbb{Z}^{D}$ or quasi-isometric to it, which seems reasonable given Figure 2) and assume the node degrees are uniformly bounded. Then, there exist positive constants $\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$ (depending on $D$ and the local geometry of $\mathcal{G}$) and an integer $h_{0}$ such that for all $h\geq h_{0}$,
$\mathfrak{C}_{1}h^{D-1}\leq|\mathcal{N}_{h}(v)|\leq\mathfrak{C}_{2}h^{D-1}.$ (6)
After having quantified the growth of the $h$-hop shell, one can prove the following theorem, which motivates our aggregation choice.
Theorem 6.2 (Dilution of mean aggregated influence in grid-like graphs).
Suppose that for a fixed $v$ and for each $h\geq h_{0}$ there exists a distinguished node $u^{*}\in\mathcal{N}_{h}(v)$ with a strong influence on $v$, quantified by $I(v,u^{*})=I^{*}>0,$ while for all other nodes $u\in\mathcal{N}_{h}(v)\setminus\{u^{*}\}$ the influence $I(v,u)$ is negligible (i.e. zero). Define $I_{\text{sum}}(v,h)=\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)=\sum_{u:\rho(v,u)=h}I(v,u)\quad\text{and}\quad I_{\text{mean}}(v,h)=\frac{1}{|\mathcal{N}_{h}(v)|}\sum_{u\in\mathcal{N}_{h}(v)}I(v,u).$ Then, $I_{\text{sum}}(v,h)\geq I^{*}$, and $I_{\text{mean}}(v,h)\leq\dfrac{I^{*}}{\mathfrak{C}_{1}\,h^{D-1}}$. Hence, as $h\to\infty$, we have $I_{\text{mean}}(v,h)\to 0,$ while $I_{\text{sum}}(v,h)$ remains bounded below by $I^{*}$. This also holds for planar graphs, where $D=2$.
Corollary 6.3 (Dilution for a planar grid-like graph).
The dilution of the mean aggregated influence for a planar grid-like graph (like our city networks) is proportional to $\frac{1}{h}$.
Corollary 6.4 (Faster dilution over aggregated $h$-hop neighborhoods).
Let $\mathcal{G}$ be a grid-like graph in $D$ dimensions, and define the aggregated $h$-hop neighborhood (or ball) of a node $v$ as $B_{h}(v)=\bigcup_{i=1}^{h}\mathcal{N}_{i}(v).$ As before, suppose that within $B_{h}(v)$ there exists a unique node $u^{*}$ with influence $I(u^{*},v)=I^{*}>0$ and that for all other nodes $u\in B_{h}(v)\setminus\{u^{*}\}$, the influence is negligible (i.e. zero). Then, the mean aggregated influence is diluted at a rate proportional to $1/h^{D}$. In particular, for a planar graph ($D=2$), the dilution occurs at a rate proportional to $1/h^{2}$.
The analysis above formally justifies the choice in Equation (2), which considers the aggregate influence of neighboring nodes $T_{h}(v)=I_{\text{sum}}$ as a more reliable measure than the mean $I_{\text{mean}}$, which is susceptible to dilution, particularly in the case of distant neighbors. This is key to the computation of the average total influence in Equation (2) and the influence-weighted receptive field in Equation (3). Finally, note the following:
Corollary 6.5 (The dilution problem does not affect the average total influence).
Let $T_{h}(v)=I_{\text{sum}}(v,h)$ be the total influence from the $h$-hop neighborhood $\mathcal{N}_{h}(v)$ of node $v$, and let $\overline{T}_{h}=\frac{1}{|V|}\sum_{v\in V}T_{h}(v)$ be the average total influence over all nodes in $V$. Suppose that for every $v$ and every $h$ there exists at least one distinguished node $u^{*}\in\mathcal{N}_{h}(v)$ satisfying $I(v,u^{*})\geq I^{*}>0.$ Then, $\overline{T}_{h}\geq I^{*},~\forall h.$
Proofs are provided in Appendix I, and a more detailed justification of Jacobian-based influence score and connections to existing literature, along with its limitations, can be found in Appendix G.
7 Conclusion
The main objective of our work is to provide better tools to help quantify long-range interactions in GNNs. Previous benchmarks, such as the LRGB (Dwivedi et al., 2023), are introduced in the context of small graph inductive learning, using solely the performance gap between classical GNNs and GTs to support the presence of long-range signals. In this work, we introduce a new large graph dataset based on city road networks, featuring long-range dependencies for transductive learning, and propose a principled measurement to quantify such dependencies. We also provide theoretical justification of both the proposed dataset and measurement, focusing on over-smoothing and influence score dilution. Beyond benchmarking purposes, our work also holds potential for a broader impact, informing applications in urban planning and transportation by providing tools to analyze and predict accessibility within city road networks.
Dataset and Code Availability
Our dataset, CityNetwork, is now publicly available in the latest version of PyTorch Geometric (PyG) as an InMemoryDataset under torch_geometric.datasets. The Total Influence metric introduced in Section 4 is also released under torch_geometric.utils. We open-sourced the code for dataset construction and benchmarking at https://github.com/LeonResearch/City-Networks, and ensure that the experimental results and figures in this work are reproducible with the provided code.
Appendix A Network Statistics
In this section, we explain the statistics used in Table 1 that characterize our dataset and discuss the estimation approach used when direct computation is impractical. We consider undirected graphs, $\mathcal{G}=(V,E)$ with $N=|V|$ nodes and $|E|$ edges, where each node $v$ is associated with a feature vector $\bm{X}_{v}\in R^{D}$ and a label $y_{v}\in\mathcal{Y}$ from a finite class set.
Node degree.
The degree $\deg(v)$ of a node $v$ represents the number of edges adjacent to it. To characterize the distribution of node degree in a network, we consider its mean $\mu_{k}$ and standard deviation $\sigma_{k}$:
$\mu_{k}=\frac{1}{N}\sum_{v\in V}\deg(v),\quad\sigma_{k}=\sqrt{\frac{1}{N}\sum_{v\in V}(\deg(v)-\mu_{k})^{2}}.$ (7)
Importantly, networks of different topologies will have different degree distributions. For example, social networks often exhibit a power-law property, where a few nodes have high degrees while most nodes have relatively low degrees. As a result, their degree distribution tends to follow a scale-free pattern with a high degree variance $\sigma_{k}^{2}$. In contrast, grid-like networks, such as city networks and super-pixel graphs, have a structured layout of connections, leading to a more uniform degree distribution with lower variance.
Clustering coefficient.
The clustering coefficient $C_{v}$ measures the tendency of a node $v\in V$ to form a tightly connected group based on triangles, and its average $\bar{C}$ measures the overall level of clustering:
$C_{v}=\frac{2T_{v}}{\deg(v)(\deg(v)-1)},\quad\bar{C}=\frac{1}{N}\sum_{v\in V}C_{v},$ (8)
where $T_{v}$ is the number of triangles that include node $v\in V$. Alternatively, transitivity offers a global measure of clustering with the following expressions:
$Transitivity=\frac{3\times\text{number of triangles}}{\text{number of connected triples}}.$ (9)
Intuitively, social networks tend to have a high average clustering coefficient and transitivity due to their community structure with highly connected hubs and frequent triadic closures. On the other hand, our city networks exhibit a low clustering coefficient and transitivity, as their structured and sparse connectivity (e.g., forming lattices) reduces the prevalence of triangles. Note that LRGB, although being “grid-like”, shows an even higher $\bar{C}$ and transitivity than social networks. This is because its networks contain many triangles, as it uses semantic super-pixels as nodes with pixel borders being the edges.
Diameter
The diameter $D$ of a network is the longest shortest path between any two nodes:
$D=\max_{u,v\in V}\rho(u,v),$ (10)
where $\rho(u,v)$ is the shortest path length from node $u$ to node $v$. Note that in this case the distance is unweighted for a more direct comparison with other datasets. It represents the maximum communication delay in the network and varies significantly across different network structures. In social networks, the presence of hubs greatly reduces the average shortest path length, leading to a small-world effect with a relatively small diameter. In contrast, grid-like networks lack hubs, and their regular structure causes the diameter to increase more rapidly as the network size grows.
Since the exact calculation of $D$ has an impractical computational complexity at $\mathcal{O}(N^{2}\log(N))$, we use the following approach to estimate the approximate diameter $\hat{D}$ of our city networks. For all nodes in a given city, we select the ones with the maximum and minimum latitude and longitude, respectively: $coord(v_{1})=(\;\cdot\;,\;\text{lat}_{max})$, $coord(v_{2})=(\;\cdot\;,\;\text{lat}_{min})$, $coord(v_{3})=(\text{long}_{max},\;\cdot\;)$, and $coord(v_{4})=(\text{long}_{min},\;\cdot\;)$. Based on these, we compute the shortest path between $(v_{1},v_{2})$ and $(v_{3},v_{4})$, and take their maximum as our final diameter estimate, that is, $\hat{D}=\max\big(d(v_{1},v_{2}),d(v_{3},v_{4})\big)$. Note that the exact diameter $D$ will always be larger than our estimation $\hat{D}$.
Homophily.
The node homophily score (Pei et al., 2020) quantifies the tendency of nodes with the same label to be connected:
$Homo=\frac{1}{N}\sum_{v\in V}\frac{|\{u\in\mathcal{N}(v)\,|\,y_{u}=y_{v}\}|}{|\mathcal{N}(v)|},$ (11)
where $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $y_{v}$ represents the label of node $v$. A higher homophily indicates a stronger preference for connections between nodes of the same class, which can significantly impact the performance of GNN models.
Appendix B Dataset Details
B.1 Raw Feature Processing
This section provides additional details of the features and labels of our City-Networks.
Node and edge features.
The node and edge features in our dataset are derived from real-world features provided by OpenStreetMap for both road junctions and road segments, as detailed below.
  • Three numerical features for the road junctions (nodes):
    • latitude: the latitude of the current road junction.
    • longitude: the longitude of the current road junction.
    • street count: the number of connected roads in both directions.
  • One categorical features for the road junctions (nodes):
    • land use: the type of land use at the current coordinate: residential, industrial, forest, farmland, commercial, railway, etc.
  • Two numerical features for road segments (edges):
    • road length: the length of the road in meters.
    • speed limit: the speed limit on the current road in km/h.
  • Two binary features for road segments (edges):
    • one-way: if the current road can only be used in one direction by vehicles.
    • reversed: if the current road alternates between different directions during rush hours in the morning and evening, which is also sometimes called "tidal flow".
  • Two categorical features for road segments (edges):
    • lanes: number of lanes in the current road, which takes either numerical values (e.g. $1,2,3,$...), or a list of numerical values (e.g. $[1,2],[2,3],[4,5],$...) when the current road has different number of lanes at different segments. We treat this feature as a categorical variable during modeling.
    • road type: the type of the current road, with possible values being: service, residential, footway, primary, secondary, tertiary, etc.
Feature engineering.
Since the categorical features land use, lanes, and road type contain many categories of only a few data points and varies across different networks, we only take the top $8$ categories with most entries for each categorical feature, and treat the rest as a single category - other. Based on our observations, we can cover more than 90% of the network with the top $8$ categories in all cases. This strategy leads to $12$ node features and $25$ edge features after one-hot encoding. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features after processing. Lastly, we transform the graph into an undirected one by merging all edges between each pair of nodes into a single edge, where the edge features are averaged using to_undirected(reduce="mean") from PyG.
B.2 Long-range Node Labels based on Local Eccentricity
Controllable long-range signal based on local eccentricity.
As mentioned in Section 2.2, we use a $k$-hop ego-network to estimate the eccentricity of each node. Specifically, after obtaining the neighborhood $\mathcal{N}_{k}(v)$ within $k$ hops of node $v$, we compute the shortest path from node $v$ to all nodes within this neighborhood, and take the maximum as the local eccentricity $\hat{\varepsilon}_{k}(v)$:
$\hat{\varepsilon}_{k}(v)=\max_{u\in\mathcal{N}_{k}(v)}\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}\sum_{e\in\pi}w(e),$ (12)
Figure 5: Distribution of the 16-hop eccentricity for all nodes in each of our City-Networks.
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ denotes the set of all possible paths from node $v$ to node $u$, and $w(e)$ represents the edge weight for edge $e$. Here, we use one of the edge features road length as the edge weight $w(e)$, such that the local eccentricity $\hat{\varepsilon}_{k}(v)$ will indicate the maximal traveling distance (in meters) from node $v$ to its $k$-hop neighbors. Figure 5 shows the distributions of such approximation at $k=16$ across different city networks, in which Paris and Shanghai have the most skewed and uniform distributions, respectively.
Importantly, this method allows us to know a priori that the long-range signal should be highly correlated with the hop $k$, which facilitates us in task design and model benchmarking. Lastly, after obtaining the local eccentricity for all nodes, we split them into $10$ quantiles as the final node labels for transductive classification.
Computational cost.
[…]
Figure 6: Visualization of local eccentricity $\hat{\varepsilon}_{k}(v)$ at $k=[8,16,32]$ and full-hop (exact eccentricity) on two sub-regions. We can observe that in the last two cases (g) and (h), node labels become highly correlated with geographic coordinates and are hence less dependent on the graph structure.
B.3 The Roles of Graph Topology and Spatial Features
The importance of structural and spatial information.
Our goal is to use a label generation approach that incorporates long-range dependency via information from both node features and graph structures. Although the local eccentricity is based on a weighted shortest-path strategy with road length being the edge weight, it does not only capture a structure perspective. This is because road length naturally relates to geographical locations, road type, land use, etc., which will link the node labels to these node features. Since the edge features are removed for the transductive task, the model will not be able to directly infer the original weighted shortest-path values. Moreover, while we can control the range of the task by changing the approximation hop $k$, this knowledge is hidden from the model. Therefore, the model needs to explore the distant neighborhood and capture both structural and spatial information to effectively infer the long-range information.
Figure 7: An example graph in PascalVOC-SP from LRGB (Dwivedi et al., 2023), where nodes are represented by super-pixels and their labels are annotated by semantics. Although spatial coordinates are highly correlated with node labels, the inductive nature makes the task still graph-dependent.
Problem of uninformative graph structures on labels.
One catastrophic problem we encountered during the experiment is that, when using the exact eccentricity or a relatively larger $k$ for node labels, the signal becomes nearly independent of graph structures. As illustrated in Figure 6, we can observe a clear correlation between the node labels and their spatial coordinates, since nodes around the city center would generally have lower eccentricity values due to the 2D grid-like topology. This means that the spatial coordinates alone are sufficient for modelling $\varepsilon(v)$, hence weakening the importance of graph structure during modeling. Similarly, some other labeling approaches would also lead to the same problem, such as using a single node as an anchor, then regarding its shortest path distance to other nodes as the node label.
Finally, we would like to point out that this problem also exists in LRGB, as illustrated by an example graph from PascalVOC-SP in Figure 7. However, while the learning task in PascalVOC-SP resembles a spatial segmentation problem on the coordinates, it still requires the model to handle graph structural information for prediction due to the inductive task nature.
Figure 8: Results of GCN at different numbers of layers on Paris with $k=[8,16,32]$ used for labels.
B.4 Choice of the Long-range Level K
As explained in Section 2.2, the choice of $k$ is critical in designing the long-range signal. On one side, $k$ should be sufficiently large to distinguish our setting from short-range datasets (e.g., social graphs with typical message passing around $4$ hops). To further elaborate on this, we follow the same experiment setups in Section 3 and test GCN on Paris using $k=[8,16,32]$ at num_layers=$[2,4,8,16,32]$, where the results are reported in Figure 8 on the right.
We can observe that, at $k=8$, the model’s performance quickly saturates as expected when num_layers reaches $8$. While at $k=16$ and $32$, the performance starts to plateau at a much deeper depth of num_layers=$16$. As such, $8$ does not seem to be an optimal choice for $k$, since there is still room to “extend” the range of the signal.
On the other hand, $k$ should also remain small enough that it does not lead to the aforementioned problem of uninformative graph structure during modelling. As visualized in Figure 6, the underlying signal, given its local nature, becomes smoother as $k$ increases from $8$ to $32$. At the same time, we can also see a increasing correlation between signal and the spatial coordinates. In the extreme case when $k$ equals the network diameter, $\hat{\varepsilon}_{k}(v)$ becomes the exact eccentricity $\hat{\varepsilon}(v)$, and the classification task resembles a “segmentation” problem on a 2D plane. This also explains the marginal performance gain from num_layers=16 to 32 at $k=32$, as the signal becomes overly smoothed that information based on $16$-hop neighborhood is sufficient to have a good “guess” of the signal at hop $32$.
At the same time, for benchmarking purposes, we wish to test GNNs and GTs with enough model depth such that they can reach the source of the signal. However, we found that when training GTs on our large city networks, due to the quadratic complexity in their global attention mechanism, they often suffer from Out-Of-Memory errors (NVIDIA-L40@48GB ) when num_layers is large.
Therefore, we decided to adopt $k=16$ in our final dataset, which we believe achieves a balance between being sufficiently long-ranged and remaining practical for computation. As discussed earlier in Section 1 and Section 2.2, we believe City-Networks brings a new and important challenge to the literature, which calls for scalable architectures that can effectively handle long-range dependency.
Appendix C Experiment Details
In this section, we first provide training details in our experiments, and then present an ablation study that supports our main claim on the long-range dependency. Lastly, we summarize the best baseline results on our dataset and discuss potential limitations of our experimental settings.
C.1 Baselines and Datasets for Comparison
Baselines.
We consider four common GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and three Graph Transformers: GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) as the baseline models in our experiments.
Datasets.
We compare the baseline results on City-Networks to the following datasets that are homophilic, heterophilic, large-scale, and long-range dependent.
  • Cora (Yang et al., 2016a) is a well-known homophilic citation network, with nodes representing documents and edges representing citation links. The features of the nodes are represented as bag-of-words that captures the content of the documents, and the goal is to predict the academic topic of each paper.
  • ogbn-arxiv (Hu et al., 2021) is a large-scale citation network of $169k$ computer science papers on arXiv that were indexed by the Microsoft academic graph, in which nodes represent papers and the directed edges indicate the citations. In particular, each node has an 128-dimensional feature vector derived from the word embeddings of titles and abstracts in the underlying papers. The task is to identify the primary category of each arXiv paper, that is, to classify each node into one of the 40 classes.
  • Amazon-Ratings (Platonov et al., 2023) is a heterophilic network that models Amazon product co-purchasing information, where nodes represent products and edges represent frequently co-purchased items. The goal is to predict average product ratings from five classes, with node features being the fastText embeddings of product descriptions.
  • PascalVOC-SP is an inductive dataset from LRGB (Dwivedi et al., 2023), which contains graphs derived from images. In particular, the nodes represent super-pixels and edges represent their boundaries. The labels are derived based on semantics, which makes the task similar to image segmentation.
As mentioned in Section 3, we closely follow the latest GNN tuning technique from Luo et al. (2024), which considers residual connection, batch normalization, dropout, etc., and use their code base for training classical GNNs and GTs on most datasets, except for PascalVOC-SP, where we adopt the hyperparameters reported by Tönshoff et al. (2024) and Shirzad et al. (2023).
C.2 Experimental Setups
As discussed earlier in Section 3, we consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all four graphs in City-Networks.
Evaluation protocols. Since our goal is to investigate the presence of long-range dependency, we train each model at num_layers=[2, 4, 8, 16] while fixing hidden_size=128, and then check if the model’s performance will positively correlate with num_layers. All cases are repeated 5 times, and we present their means and standard deviations. In the ablation studies below, we also investigate different choices of hidden_size in [16, 32, 64, 128] following this strategy, which not only acts as hyperparameter tuning, but also shows the robustness of our conclusion to different choices of model hyperparameters.
Hyperparameters
We run each model for $30k$ maximum epochs using AdamW (Loshchilov & Hutter, 2019) optimizer for sufficient training, during which we record the validation accuracy every $100$ epochs and save the model at the best validation epoch for final testing. Meanwhile, we closely follow the latest GNN tuning technique from Luo et al. (2024), and our hyperparameter search space is summarized in Table 4. Note that for GTs, we do not apply positional encoding due to its impractical computation on our large graphs (which is one of the challenges of applying GTs on large graphs), and use GCN as their internal MPNNs.
Table 4: Summary of hyperparameters and their search space.
Type Hyper-parameter Search range Default
Model num_layers [2, 4, 8, 16][2,\ 4,\ 8,\ 16] $16$
hidden_size [16, 32, 64, 128][16,\ 32,\ 64,\ 128] $128$
pre_linear_layer [0, 1, 2][0,\ 1,\ 2] 0
post_linear_layer [0, 1, 2][0,\ 1,\ 2] 22
residual [True, False] True
Train learning rate $10^{-4},\ 5\times 10^{-4},\ 10^{-3}]$ $10^{-3}$
dropout [0, 0.2, 0.5, 0.7][0,\ 0.2,\ 0.5,\ 0.7] $0.2$
weight decay $0,\ 10^{-5},\ 5\times 10^{-5},\ 10^{-4}]$ $10^{-5}$
normalization [None, BatchNorm, LayerNorm] BatchNorm
C.3 Ablation Studies
Results under different hidden channel sizes.
To further support our main findings, we test two classical GNN baselines: GCN (Kipf & Welling, 2017) and GraphSAGE (Hamilton et al., 2017); and two GT baselines: Exphormer (Shirzad et al., 2023) and SGFormer (Wu et al., 2023) under different hidden_size=[16, 32, 64, 128] on Paris, Shanghai, and L.A.. The results are presented in Figure 9, where we can observe a consistent pattern with our main findings in Figure 3—the baselines’ performance improves substantially when increasing the number of layers from $2$ to $16$.
Notably, all baselines generally achieve superior performance with a larger hidden_size. However, such differences are often negligible compared to the increasing trend in model depths. With that being said, in the next section, we proceed to investigate scenarios when the model size is fixed.
Results at fix model depth with various sampling hops.
As discussed earlier, one limitation of the main results on Figure 3 is that the observed performance gains may be attributed to the increasing model parameters as the number of layers $L$ grows. To address this, we further investigate the scenario where $L$ is fixed at 16 (i.e., keeping the model size constant), and adopt an $H$-hop neighborhood sampling method introduced in GraphSAGE (Hamilton et al., 2017) that refrains the model from seeing beyond hop $H$. Note that the same method is also widely used in the literature for training models on large graphs.
Concretely, our strategy is implemented with NeighborLoader from PyG, which recursively selects $N_{1},N_{2},\dots,N_{H}]$ neighbors from a node’s 1st, 2nd, …, $H$-th hop neighborhood. Given the grid-like structure of our city networks, the size of the neighborhood does not explode with $H$, and we empirically found that the average size of a $16$-hop ego-network is typically around $1k$ nodes, which remains manageable for most GPU devices. Therefore, we sample all nodes inside the $H$-hop neighborhood, and use a batch_size of $20k$ (i.e., $20k$ seed nodes) for training and testing.
The results are presented in Figure 10, where we can observe that every baseline, at a fixed model architecture, shows a consistent upward trend on all four city networks. These results indicate that long-range dependency, rather than the parameter size, is the primary factor that contributes to the improvement of the model’s performance.
One may notice the trend from $H=8$ to $H=16$ remains relatively flat. We attribute this phenomenon to the overlapped subgraphs sampled from NeighborLoader, where the ego-networks of different seed nodes often share common supporting nodes. Consequently, if the number of layers $L$ exceeds $H$, the model can capture information from more distant hops beyond $H$, rather than being limited to a maximum of $H$ as expected. This effect is further amplified when using a large batch size (i.e., a large number of seed nodes) or a large number of hops (i.e. a large neighborhood size), as it increases the chance of overlapping ego-networks. As a result, the performance remains similar between these two settings. Note that this phenomenon is not a contradiction to our main claims, but rather a limitation, as we can not create perfect 8-hop subgraphs.
Figure 9: Ablation results under hidden_size=[16, 32, 64, 128] and num_layers=[2, 4, 8, 16] on Paris, Shanghai, and Los Angeles. The patterns are all consistent with our findings in Section 3.
Figure 10: Results for fixing number of layers $L=16$ and setting numbers of hops $H\in[2,4,8,16]$.
C.4 Summary of Baseline Results
For reference, we summarize the best baseline results on our City-Networks in Table 5. Note that due to the global attention module in GraphGPS, it is highly memory-consuming and fails to fit into our GPU memory (4848GB), even with hidden_size=$16$ under num_layers=$16$.
Table 5: Baseline results on City-Networks.
Baseline Type Paris Shanghai Los Angeles London
MPNNs MLP $25.5\pm 0.4$ $28.4\pm 0.6$ $24.1\pm 0.5$ $27.9\pm 0.1$
ChebNet $54.1\pm 0.2$ $66.5\pm 0.1$ $61.4\pm 0.4$ $54.7\pm 0.2$
GCN $53.2\pm 0.3$ $62.1\pm 0.2$ $58.3\pm 0.3$ $50.1\pm 0.7$
GraphSAGE $54.6\pm 0.2$ $68.3\pm 0.5$ $61.4\pm 0.3$ $55.4\pm 0.2$
GAT $51.1\pm 0.3$ $68.0\pm 0.5$ $59.5\pm 0.3$ $52.0\pm 0.3$
GCNII $51.3\pm 0.2$ $61.5\pm 0.4$ $56.0\pm 0.3$ $48.2\pm 0.3$
DropEdge $48.2\pm 0.2$ $60.8\pm 0.4$ $55.5\pm 0.3$ $45.0\pm 0.3$
GTs GraphGPS $52.1\pm 0.6$ $63.0\pm 0.5$ $59.8\pm 0.5$ OOM
Exphormer $55.1\pm 0.8$ $70.2\pm 0.4$ $63.8\pm 0.6$ $49.5\pm 0.4$
SGFormer $52.0\pm 0.8$ $64.1\pm 0.3$ $60.1\pm 0.7$ $48.3\pm 0.3$
Appendix D Computational Complexity in R
The computation of $R$ requires calculating $T_{h}(v)$ for $v\in V$ at $h\in[0,1,...,H]$, in which the dominant cost stems from computing the Jacobian matrix. Since the model’s gradient at node $v$ will be zero for nodes beyond its $H^{th}$ hop, we only need to compute the Jacobian within each node’s $H$-hop neighborhood. This leads to a computational cost of $\mathcal{O}(N\bar{N}_{H})$, with $\bar{N}_{H}$ being the average size of $H$-hop ego-networks on $\mathcal{G}$:
$\bar{N}_{H}=\frac{1}{N}\sum_{v\in V}|\mathcal{N}_{H}(v)|,\quad\mathcal{N}_{H}(v)=\{u\in V\,|\,d(v,u)\leq H\},$ (13)
where $d(v,u)$ is the length of shortest path from node $v$ to node $u$. Given the large scale of our city networks, we employ a stochastic approximation that samples $10k$ nodes to further reduce the computational cost. For reference, the calculation finishes under 30 minutes for all baselines on all four City-Networks using a single NVIDIA RTX 3090 GPU with 48 AMD Ryzen 3960XCPU cores.
Appendix E Limitations in Experiments
Total Influence on large dense graphs.
During our experiment, we found that on large dense graphs, especially the one with small diameters such as ogbn-arxiv ($|V|=169k$, $diam=25$), it is difficult to compute our Total Influence measurement, as $\bar{N}_{H}$ in Equation (13) quickly converges to $N$ as $H$ increases. In this case, if the model also happens to have complex architectures (e.g., GraphGPS), the computation of $R$ and $T_{h}$ will therefore become impractical. We believe this constitutes a limitation of our proposed measurement, where future works may explore more effective methods for dense graphs based on stochastic sampling.
Appendix F Proofs and Justifications in Section 5
F.1 Details of Linearization
A Graph Convolutional Network (GCN) layer with input features $\bm{X}^{(l)}$, learnable weights $\bm{W}^{(l)}$ for layer $l$, and non-linear pointwise activation function $\sigma$ is defined as:
$\bm{X}^{(l+1)}=\sigma\left(\bm{\tilde{S}}_{adj}\bm{X}^{(l)}\bm{W}^{(l)}\right).$ (14)
As pointed out in Wu et al. (2019), under linearization (removal of non-linear activations), an $L$-layer Simple Graph Convolution (SGC) can be expressed as:
$\bm{X}^{(L)}=(\bm{\tilde{S}}_{adj})^{L}\bm{X}^{(0)}(\prod_{l=0}^{L-1}\bm{W}^{(l)}),$ (15)
where $(\bm{\tilde{S}}_{adj})^{L}$ is the $L$-th power of the normalized adjacency operator. The result follows directly from iterative application of the linearized GCN operation.
Next, it is well known that the normalized adjacency operator $\bm{\tilde{S}}_{adj}$ admits an eigendecomposition:
$\bm{\tilde{S}}_{adj}=\bm{\tilde{U}}\bm{\Lambda}\bm{\tilde{U}}^{T},$ (16)
where $\bm{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}<\ldots<\lambda_{N}$ and $\bm{\tilde{U}}$ contains the corresponding eigenvectors. And, for any positive integer $k$, the power iteration is as follows:
$(\bm{\tilde{S}}_{adj})^{k}=\bm{\tilde{U}}\bm{\Lambda}^{k}\bm{\tilde{U}}^{T},$ (17)
where $\bm{\Lambda}^{k}=\text{diag}(\lambda_{1}^{k},\ldots,\lambda_{N}^{k})$.
Wu et al. (2019) discuss the spectral properties of the normalized adjacency operator, which satisfies: $\lambda_{N}=1$, and $|\lambda_{i}|<1$ for all $i<N$. Therefore, as the number of layers approaches infinity, the layer collapse phenomenon is well-documented in the literature:
$\lim_{l\to\infty}(\bm{\tilde{S}}_{adj})^{l}=\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T},$ (18)
where $\bm{\tilde{u}}_{n}$ is the eigenvector corresponding to $\lambda_{N}$. This is because as $l\to\infty$, $\lambda_{i}^{l}\to 0$ for all $i<N$ since $|\lambda_{i}|<1$. Meanwhile, $\lambda_{N}^{l}=1^{l}=1$ for all $l$. Thus,
$\lim_{l\to\infty}(\bm{\tilde{S}}_{adj})^{l}=\bm{\tilde{U}}\begin{pmatrix}0&&\\ &\ddots&\\ &&1\end{pmatrix}\bm{\tilde{U}}^{T}=\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}.$ (19)
This ultimately leads to information loss. In the limit as $k\to\infty$, the graph convolution operation collapses all node features to scalar multiples of $\bm{\tilde{u}}_{n}$, the entry of which at node $v$ is $\sqrt{1+\text{degree}(v)}$, resulting in complete loss of the original feature information. In other words, the learned representations suffer from over-smoothing.
F.2 Proofs for Theoretical Results in Section 5
Lemma F.1 (Eigenvalue complementarity of normalized operators).
For a connected graph, the eigenvalues $\bm{{S}}_{adj}$ and $\bm{L}_{sym}$ exhibit the following complementarity relationship:
$\lambda_{N+1-i}(\bm{{S}}_{adj})=1-\lambda_{i}(\bm{L}_{sym})$ (20)
for all $i=1,\ldots,N$, where $N$ is the number of nodes in the graph, and eigenvalues (for both operators) are indexed such that $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$.
Proof of Lemma 20.
By definition, $\bm{L}_{sym}=\bm{I}-\bm{S}_{adj}$. Since both matrices are symmetric, they are diagonalizable with real eigenvalues. Let $\bm{u}$ be an eigenvector of $\bm{L}_{sym}$ with eigenvalue $\lambda_{i}(\bm{L}_{sym})$. Then:
$\bm{L}_{sym}\bm{u}=\lambda_{i}(\bm{L}_{sym})\bm{u}\Rightarrow(\bm{I}-\bm{S}_{adj})\bm{u}=\lambda_{i}(\bm{L}_{sym})\bm{u}\Rightarrow\bm{S}_{adj}\bm{u}=(1-\lambda_{i}(\bm{L}_{sym}))\bm{u}.$ (21)
Therefore, $\bm{u}$ is also an eigenvector of $\bm{S}_{adj}$ with eigenvalue $1-\lambda_{i}(\bm{L}_{sym})$. Since the eigenvalues are ordered in ascending order for $\bm{L}_{sym}$ and the transformation $1-\lambda_{i}$ reverses this ordering, we have $\lambda_{N+1-i}(\bm{{S}}_{adj})=1-\lambda_{i}(\bm{L}_{sym})$. ∎
This complementarity implies that when the normalized algebraic connectivity $\lambda_{2}(\bm{L}_{sym})$ is small, the second largest positive eigenvalue of $\bm{{S}}_{adj}$ must be close to 1. However, in our context, we are interested in the algebraic connectivity of the normalized graph Laplacian that would correspond to the normalized adjacency operator, $\bm{\tilde{S}}_{adj}$, introduced in Section 5.1, instead of that of $\bm{{S}}_{adj}$, motivating the following:
Proposition F.2 (Self-loops decrease algebraic connectivity of the original graph, from Section 5.2).
Assume a connected graph $\mathcal{G}$ with more than two nodes. For all $\gamma>0$,
$\lambda_{N-1}\left(\bm{S}_{adj}\right)=\lambda_{N-1}\left(\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}\right)<\lambda_{N-1}\left((\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}(\gamma\bm{I}+\bm{A})(\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}\right).$ (22)
Proof of Proposition 4 (same as Proposition 22).
Let $\mathcal{G}^{\prime}$ be the graph $\mathcal{G}$ with self-loops added, each having weight $\gamma>0$. If $\mathcal{G}$ already contains self-loops, their weights are increased by $\gamma$. We denote the vertex set of $\mathcal{G}^{\prime}$ as $V$, with the obvious correspondence to the vertices of $\mathcal{G}$. Then $\bm{L}_{sym}^{\mathcal{G}}=\bm{I}-\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}$ and $\bm{L}_{sym}^{\mathcal{G}^{\prime}}=\bm{I}-(\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}(\gamma\bm{I}+\bm{A})(\gamma\bm{I}+\bm{D})^{-\frac{1}{2}}$, so proving
$\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}})>\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}^{\prime}}).$ (23)
will prove the proposition. Note that in practice we care about the case where $\gamma=1$. We proceed as follows: we take the eigenfunction on $\mathcal{G}$ corresponding to $\lambda_{\mathcal{G}}$, lift it to $\mathcal{G}^{\prime}$, and show that this yields an upper bound. Using the variational characterisation of eigenvalues in (Chung, 1997, Eq. 1.13) and the fact that edge weights $w_{\mathcal{G}}(u,v)=w_{\mathcal{G}^{\prime}}(u,v)$ if $u\neq v$, and the degrees of vertices $d^{\mathcal{G}^{\prime}}_{v}=\gamma+d^{\mathcal{G}}_{v}$:
$\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}})$ $=\inf_{f\,:\,\sum_{x\in V}f(x)d^{\mathcal{G}}_{x}=0}\frac{\sum_{x\sim y}(f(x)-f(y))^{2}w(x,y)}{\sum_{x\in V}f(x)^{2}d^{\mathcal{G}}_{x}}$ (24)
Pick such an $f$ attaining this infimum (i.e., the eigenvector of $\bm{L}_{sym}^{\mathcal{G}}$ corresponding to $\lambda_{2}(\bm{L}^{\mathcal{G}}_{sym})$ multiplied by $\bm{D}^{-\frac{1}{2}}$). We use this $f$ to construct a signal $g$ on $\mathcal{G}^{\prime}$. Let
$g(x)=f(x)-\frac{\gamma\sum_{x\in V}f(x)}{\sum_{x\in V}d^{\mathcal{G}}_{x}+\gamma}.$ (25)
That is, we reduce $f$ by a constant everywhere. We have picked the constant such that
$\sum_{x\in V}g(x)d^{\mathcal{G}^{\prime}}_{x}=\sum_{x\in V}g(x)(d^{\mathcal{G}}_{x}+\gamma)=0.$ (26)
Since $g$ is simply $f$ shifted by a constant,
$\forall x,y\in V:\quad g(x)-g(y)=f(x)-f(y).$ (27)
Again, by the variational characterisation of eigenvalues in (Chung, 1997, Eq. 1.13):
$\lambda_{2}(\bm{L}_{sym}^{\mathcal{G^{\prime}}})$ $=\inf_{h\,:\,\sum_{x\in V}h(x)d^{\mathcal{G}^{\prime}}_{x}=0}\frac{\sum_{x\sim y}(h(x)-h(y))^{2}w(x,y)}{\sum_{x\in V}h(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}$ (28)
$\leq\frac{\sum_{x\sim y}(g(x)-g(y))^{2}w(x,y)}{\sum_{x\in V}g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}=\frac{\sum_{x\sim y}(f(x)-f(y))^{2}w(x,y)}{\sum_{x\in V}g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}$ (29)
The first inequality holds because $g$ is in the set $\{{h\,:\,\sum_{x\in V}h(x)d^{\mathcal{G}^{\prime}}_{x}=0}\}$ , and the infimum serves as a lower bound for the function on any element of that set. The last equality follows from applying (27).
We now show that ${\sum_{x\in V}g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}>{\sum_{x\in V}f(x)^{2}d^{\mathcal{G}}_{x}}$. This will let us bound (29) above by (24). Expanding the definition of $g$:
$\sum g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}$ $=\sum f(x)^{2}d^{\mathcal{G^{\prime}}}_{x}-2\frac{\gamma(\sum f(x)d^{\mathcal{G^{\prime}}}_{x})(\sum f(x))}{\sum d^{\mathcal{G^{\prime}}}_{x}}+\frac{\gamma^{2}(\sum f(x))^{2}}{\sum d^{\mathcal{G^{\prime}}}_{x}}$ (30)
Noting that $\sum f(x)d^{\mathcal{G}}_{x}=0$ and $d^{\mathcal{G^{\prime}}}_{x}=d^{\mathcal{G}}_{x}+\gamma$ so $\sum f(x)d^{\mathcal{G^{\prime}}}_{x}=\gamma\sum f(x)$,
$\sum g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}-\sum f(x)^{2}d^{\mathcal{G}}_{x}$ $=\gamma\sum f(x)^{2}-2\frac{\gamma^{2}(\sum f(x))^{2}}{\sum d^{\mathcal{G^{\prime}}}_{x}}+\frac{\gamma^{2}(\sum f(x))^{2}}{\sum d^{\mathcal{G^{\prime}}}_{x}}$ (31)
$=\gamma\sum f(x)^{2}-\gamma^{2}\frac{(\sum f(x))^{2}}{\sum d^{\mathcal{G^{\prime}}}_{x}}$ (32)
By the Cauchy-Schwarz inequality on $f(\bm{x})$ and $\bm{1}$,
$(\sum f(x))^{2}=(\sum f(x)\cdot 1)^{2}\leq(\sum f(x)^{2})(\sum 1^{2})=n\sum f(x)^{2}.$ (33)
Furthermore, as $\mathcal{G}$ is connected, each node has degree of at least 1 – $\forall x:d^{\mathcal{G}}_{x}\geq 1$. As the graph has more than two nodes, one node must have degree of at least 2. So $\sum d^{\mathcal{G^{\prime}}}_{x}>n(1+\gamma)$, and therefore:
$\sum g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}-\sum f(x)^{2}d^{\mathcal{G}}_{x}$ $\geq\frac{\gamma}{n}(\sum f(x))^{2}-\gamma^{2}\frac{(\sum f(x))^{2}}{n(1+\gamma)}$ (34)
$=\frac{(\sum f(x))^{2}}{n}\left(\gamma-\frac{\gamma^{2}}{1+\gamma}\right)$ (35)
$=\frac{(\sum f(x))^{2}}{n}\left(\frac{\gamma}{1+\gamma}\right)$ (36)
$>0.$ (37)
Based on which we conclude that ${\sum_{x\in V}g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}>{\sum_{x\in V}f(x)^{2}d^{\mathcal{G}}_{x}}$, and so by (24) and (29):
$\lambda_{2}(\bm{L}_{sym}^{\mathcal{G^{\prime}}})$ $\leq\frac{\sum_{x\sim y}(f(x)-f(y))^{2}w(x,y)}{\sum_{x\in V}g(x)^{2}d^{\mathcal{G^{\prime}}}_{x}}$ (38)
$<\frac{\sum_{x\sim y}(f(x)-f(y))^{2}w(x,y)}{\sum_{x\in V}f(x)^{2}d^{\mathcal{G}}_{x}}=\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}}).$ (39)
Hence, $\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}})>\lambda_{2}(\bm{L}_{sym}^{\mathcal{G}^{\prime}})$ and by (23), the proof is complete. ∎
Proof of Theorem 5.
Given Lemma 20 (correspondence between the eigenvalues of $\bm{S}_{adj}$ and $\bm{L}^{\mathcal{G}}_{sym}$) and Proposition 4, seeing that $diam(\mathcal{G})\geq 4$ means that the graph has more than two nodes, the result follows immediately from (Chung, 1997, Lemma 1.14). ∎
F.3 Rationale in Theorem 5
The second largest eigenvalue of $\bm{\tilde{S}}_{adj}$ in terms of magnitude is either $\lambda_{N-1}$ or $\lambda_{1}$ . We explicitly consider the case when it is $\lambda_{N-1}$. As $\lambda_{N-1}$ approaches 1, the rate of convergence to the limiting state $\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}$ decreases exponentially with the number of layers. For any layer $l$, the difference from the limiting state can be expressed as:
[…]
where $||\cdot||$ denotes the spectral norm and the last equality follows from that the spectral norm of a diagonal matrix being equal to the largest absolute value of its diagonal entries. Thus, when $\lambda_{N-1}$ is close to 1, more layers are required to achieve the same level of convergence to the limiting state. Since $\lambda_{N-1}<1$, taking powers of $\lambda_{N-1}$ will eventually converge to zero, but the rate of this convergence slows dramatically as $\lambda_{N-1}$ approaches 1. Hence, because graphs with a large diameter and low maximum degree have a lower algebraic connectivity due to reduced inter-component connectivity, GNNs operating on such graphs will be less susceptible to over-smoothing when processing node features.
We have focused on the case where the second largest eigenvalue by magnitude of $\bm{\tilde{S}}_{adj}$ is $\lambda_{N-1}$. The special case where it is instead $\lambda_{1}$ (which, in this case, must be negative) gives a different regime, where it is possible for graphs with smaller diameters to exhibit less over-smoothing than those with larger diameters. Consider, for example, a bipartite graph where over-smoothing occurs independently on each side of the partition, resulting in two distinct values depending on which partition the node is in, rather than convergence to a multiple of $\bm{\tilde{u}}_{N}$, where the node feature value would only depend on the degree.
Nevertheless, our analysis demonstrates that large diameters and sparse connectivity generally reduce the likelihood of over-smoothing. This insight motivates our proposal of benchmark datasets with significantly larger diameters than those currently used in the literature (see Table 1). The underlying hypothesis is that reduced likelihood of over-smoothing in high-diameter, sparse networks enables the possibility of GNNs learning representations that capture long-range dependencies when necessary.
Appendix G Jacobian-based Influence Score Motivation
Here, we provide additional motivation behind the proposed measurement for quantifying long-rangeness in Section 4.
G.1 Intuition behind Jacobian-based Influence Score
The Jacobian has been used in analysis of node interactions in GNNs in multiple previous works (Xu et al., 2018; Gasteiger et al., 2022; Di Giovanni et al., 2023). For example, it is used in Di Giovanni et al. (2023, Theorem 4.1) to show when over-squashing happens in long-range interactions, and to show how vanishing gradients occur in very deep GNNs. Influence specifically has been used to compute a natural measure of interactions between two nodes (Xu et al., 2018). We accordingly use aggregated influence, Equation (2), to gauge how nodes at a distance $h$ affect the output of the GNN at a focal node, thus quantifying long-rangeness. By the definition of partial derivatives, we can understand the Jacobian as follows:
$\partialderivative{\bm{H}^{(\ell)}_{vi}(\bm{X})}{\bm{X}_{uj}}=\lim_{\delta\to 0}\frac{\bm{H}^{(\ell)}_{vi}(\bm{X}+\delta\bm{e}_{uj})-\bm{H}^{(\ell)}_{vi}(\bm{X})}{\delta},$ (41)
where $\bm{X}$ is the original (unperturbed) input feature matrix, and $\delta\bm{e}_{uj}$ is an infinitesimal perturbation in the $j^{th}$ component of the feature vector at node $u$ (a standard basis vector in the node feature space). This means the more positive the Jacobian entry is, the more a positive perturbation of the features at node $u$ and component $j$ will increase the logits at node $v$ and component $i$ at the final layer. In other words, the Jacobian entry being positive or negative means that the logits are pushed up or down. Given that, after applying the softmax function, the probabilities at a given data point increase monotonically with the logits at that point, we can consider both positive and negative influences as actual influence and only focus on the absolute value (i.e., sensitivity rather than direction).
Monotonicity of the Softmax function.
Consider the derivative of the $i$-th softmax probability with respect to its corresponding logit:
$\partialderivative{p_{i}}{z_{i}}=\partialderivative{z_{i}}\left(\frac{\exp(z_{i})}{\sum_{j}\exp(z_{j})}\right)=\frac{\exp(z_{i})\sum_{j}\exp(z_{j})-\exp(z_{i})\exp(z_{i})}{(\sum_{j}\exp(z_{j}))^{2}}=p_{i}(1-p_{i})$ (42)
Observe that: $p_{i}(1-p_{i})>0$ for $0<p_{i}<1$ and $\partialderivative{p_{i}}{z_{i}}\to 0$ as $p_{i}\to 0$ or $p_{i}\to 1$. For $j\neq i$:
$\partialderivative{p_{i}}{z_{j}}=\partialderivative{z_{j}}\left(\frac{\exp(z_{i})}{\sum_{h}\exp(z_{h})}\right)=-\frac{\exp(z_{i})\exp(z_{j})}{(\sum_{h}\exp(z_{h}))^{2}}=-p_{i}p_{j}$ (43)
This proves that the softmax probabilities increase with their corresponding logits and decrease with other logits. ∎
The above is a well-known fact and we do not claim novelty.
G.2 Potential Limitations: Output Cancellation
However, it is possible to find counterexamples in which measuring the absolute Jacobian sensitivity could be insufficient or even misleading: if a positive and negative influence always cancel each other out. We find such a situation can happen in unregularised linear models with heavy collinearity of features (Hastie et al., 2009, p.63) – indeed, this is presented as one of the motivations of ridge regression. To better understand this, we give a simple model of such cancellation:
Proposition G.1 (Absolute Jacobian sensitivity may over-estimate influence).
There exists a model $h_{v}$, where the combined effect of changes in input variables on the output is zero (i.e., the sum of the partial derivatives is zero), while the sum of the absolute values of the individual partial derivatives is nonzero.
Proof of Proposition G.1.
Consider a graph with three nodes $u,v,w$, where $v$ is the focal node for our calculation. Let input and output features be scalars $\mathbf{X}_{u}=x_{u}$ on $\mathbf{H}_{v}=h_{v}$ (nodewise binary classification where using a single logit as input to a sigmoid function is possible). The influence is $I(v,u)=\left|\frac{\partial h_{v}}{\partial x_{u}}\right|$, and similarly for $\mathbf{X}_{w}=x_{w}$, $I(v,w)=\left|\frac{\partial h_{v}}{\partial x_{w}}\right|$. Assume the model learns the function $h_{v}=x_{u}-x_{w}$ and that $x_{u}\approx x_{w}$, then following the definition in the main text:
$T_{h}(v)=I(v,u)+I(v,w)=\left|\frac{\partial h_{v}}{\partial x_{u}}\right|+\left|\frac{\partial h_{v}}{\partial x_{w}}\right|=|1|+|-1|=2,$ (44)
while $h_{v}=0.$ In fact,
$\frac{\partial h_{v}}{\partial x_{u}}+\frac{\partial h_{v}}{\partial x_{w}}=1+(-1)=0$ (45)
holds for all $(x_{u},x_{w})$; no small-difference assumption on $x_{u}$ and $x_{w}$ is required.
This demonstrates that the sum of the absolute values of the Jacobian entries can be non-zero, even when the net effect on the output is zero. The key is the opposing nature of their influence, not their specific values. The example sets $x_{u}$ and $x_{w}$ as approximately equal to highlight that the net output can be small (or changes to it can be small) while the individual influences are significant. ∎
Although this is a valid concern, we next show that for Message Passing Neural Networks (MPNNs), at least at initialization, such cancellation does not happen.
Definition G.2 (Message-Passing Neural Network layer).
For a MPNN layer $l$, the node feature update for $v$ is given by: $\bm{X}_{v}^{(l+1)}=\phi\left(\bm{X}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}\psi\left(\bm{X}_{v}^{(l)},\bm{X}_{u}^{(l)}\right)\right),$ where $\psi$ is a message function, responsible for computing interactions between neighboring nodes, $\bigoplus$ is a permutation-invariant aggregation function, such as summation, mean, or max, $\phi$ is an update function that integrates aggregated information into the node representation.
Definition G.3 (Smooth Hypersurface in $\mathbb{R}^{D}$).
A hypersurface in $\mathbb{R}^{D}$ is a subset defined locally as the zero set of a continuously differentiable function $f:\mathbb{R}^{D}\to\mathbb{R}$ such that the gradient $\nabla f$ is nonzero at almost every point where $f=0$. That is, if we have an equation of the form $f(\bm{\theta})=0$, where $\bm{\theta}$ is a vector of parameters, and if $\nabla f(\bm{\theta})\neq 0$ generically, then $f=0$ defines a (locally) (D1)(D-1)-dimensional manifold, which is a hypersurface.
Theorem G.4 (Measure-zero of exact cancellation at MPNN initialization).
Consider an MPNN where the functions $\psi$ and $\phi$ are parameterized by $\bm{\theta}$ and differentiable, typically modeled as MultiLayer Perceptrons (MLPs). Suppose the parameters $\bm{\theta}$ are drawn from a probability distribution that is absolutely continuous with respect to the Lebesgue measure. Then, the set of parameter configurations for which exact cancellation of Jacobian contributions occurs has Lebesgue measure zero.
Proof of Theorem G.4.
For exact cancellation to hold assuming $\bigoplus=\sum$, the following sum must be identically zero while the individual terms remain non-zero. Since $\psi$ and $\phi$ are differentiable and parameterized by $\bm{\theta}$, each Jacobian term is a smooth function of $\bm{\theta}$. The equation:
$f(\bm{\theta})=\sum_{u\in\mathcal{N}(v)}\frac{\partial\psi(\bm{X}_{v}^{(l)},\bm{X}_{u}^{(l)})}{\partial\bm{X}_{u}^{(l)}}=0,$ (46)
defines a level set of smooth functions, a hypersurface (or a set of lower-dimensional submanifolds) in parameter space. The solution set of a nontrivial smooth function has Lebesgue measure zero unless it is identically zero across all $\bm{\theta}$, which is not the case here. Furthermore, since $\bm{\theta}$ is drawn from an absolutely continuous distribution (such as Gaussian or uniform), the probability of exactly selecting a parameter that lies on this hypersurface is zero at initialization. Thus, exact cancellation of Jacobian contributions is an event of measure zero in the space of our idealized single-layer MPNN: this can naturally be extended to multiple layers. Note that in this proof we have assumed the Jacobian sum is not identically zero by construction, unlike in Proposition G.1 where the function was explicitly constructed to ensure cancellation. ∎
Lastly, it is worth noting an alternative perspective on interpreting influence measures derived from the sum of absolute Jacobian entries, particularly when considering potential cancellation effects as demonstrated in Proposition G.1. One can indeed argue that the measure’s utility lies precisely in its capacity to quantify the magnitude of sensitivity to inputs from individual distant nodes or pathways, irrespective of whether these influences ultimately negate one another in contributing to the final output. From this stance, focusing on the sum of absolute values reveals the underlying structure and strength of potential long-range dependencies (the information channels themselves) that might be obscured by observing only the net effect. This interpretation, therefore, hinges on defining dependency or interaction based on the existence and intensity of these information flow pathways, rather than strictly on their final, combined impact on a node’s prediction.
Appendix H Additional Definitions in Section 6
Definition H.1 (Standard lattice in $\mathbb{R}^{D}$).
The standard lattice in $\mathbb{R}^{D}$, denoted by $\mathbb{Z}^{D}$, is the set of all integer-coordinate points in $\mathbb{R}^{D}$: $\mathbb{Z}^{D}=\{(z_{1},z_{2},\dots,z_{D})\mid z_{i}\in\mathbb{Z}\text{ for all }i=1,\dots,D\}.$ Equivalently, $\mathbb{Z}^{D}$ consists of all points that can be written as integer linear combinations of the standard basis vectors: $\mathbb{Z}^{D}=\left\{\sum_{i=1}^{D}z_{i}\bm{e}_{i}\mid z_{i}\in\mathbb{Z}\right\},$ where $\{\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{D}\}$ is the standard basis for $\mathbb{R}^{D}$, meaning each $\bm{e}_{i}$ is a unit vector with a 1 in the $i$-th position and 0 elsewhere. This lattice forms a grid-like structure in $\mathbb{R}^{D}$ with each point having exactly $2D$ one-hop (adjacent lattice) neighbors. For a planar graph $D=2$, hence, each node has a total of 4 neighbors.
Definition H.2 ($h$-hop shells).
Let $\mathcal{G}=(V,E)$ be a graph with shortest-path distance $\rho:V\times V\to\mathbb{N}$. The $h$-hop shell (or $h$-hop neighborhood) of a node $v\in V$ is defined as $\mathcal{N}_{h}(v)=\{u\in V:\rho(v,u)=h\}.$ That is, $\mathcal{N}_{h}(v)$ consists of all nodes that are exactly $h$ hops away from $v$.
Definition H.3 (Quasi-isometric graphs).
Let $\mathcal{G}_{1}=(V_{1},E_{1})$ and $\mathcal{G}_{2}=(V_{2},E_{2})$ be two graphs equipped with shortest-path distance functions $\rho_{1}:V_{1}\times V_{1}\to\mathbb{R}_{\geq 0}$ and $\rho_{2}:V_{2}\times V_{2}\to\mathbb{R}_{\geq 0}$, respectively (in our case distances are in $\mathbb{N}$). We say that $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are quasi-isometric if there exist constants $\lambda\geq 1$, $\mathfrak{C}\geq 0$, and $\mathfrak{D}\geq 0$, and a function $f:V_{1}\to V_{2}$ such that for all $u,v\in V_{1}$: $\frac{1}{\lambda}\rho_{1}(u,v)-\mathfrak{C}\leq\rho_{2}(f(u),f(v))\leq\lambda\rho_{1}(u,v)+\mathfrak{C}$, Every node in $V_{2}$ is within distance $\mathfrak{D}$ of some $f(u)$, i.e., $\forall v\in V_{2},\exists u\in V_{1}\text{ such that }\rho_{2}(v,f(u))\leq\mathfrak{D}.$
Appendix I Proofs for Theoretical Results in Section 6
Proof of Lemma 6.
Since $\mathcal{G}$ is grid-like in $D$ dimensions, its structure mimics that of $\mathbb{Z}^{D}$. In $\mathbb{R}^{D}$, the volume of a ball of radius $h$ scales as $h^{D}$ (recall the volume of a sphere in $\mathbb{R}^{3}$ is $\frac{4}{3}\pi h^{3}$). Let
$B_{h}(v)=\{u\in V:\rho(v,u)\leq h\},$ (47)
denote the ball of radius $h$ centered at $v$. Then for large $h$ (asymptotic bound),
$|B_{h}(v)|=\Theta(h^{D}).$ (48)
Since the $h$-hop neighborhood is the set difference
$\mathcal{N}_{h}(v)=B_{h}(v)\setminus B_{h-1}(v).$ (49)
This effectively means that the size of $\mathcal{N}_{h}(v)$ is approximately the difference between the volume of two consecutive balls:
$|\mathcal{N}_{h}(v)|=|B_{h}(v)|-|B_{h-1}(v)|.$ (50)
A standard asymptotic argument implies
$|\mathcal{N}_{h}(v)|=\Theta(h^{D-1}),$ (51)
since we can approximate the aforementioned difference via:
$|B_{h}(v)|-|B_{h-1}(v)|=\mathfrak{C}h^{D}-\mathfrak{C}(h-1)^{D}\approx\mathfrak{C}h^{D}-\mathfrak{C}(h^{D}-Dh^{D-1})=\mathfrak{C}Dh^{D-1},$ (52)
using the binomial expansion and assuming large $h$ (we are concerned with long-range interactions), where the leading order term dominates. Intuitively, this corresponds to the surface growth of the ball, which in $\mathbb{R}^{D}$ scales as $h^{D-1}$ (the area of a sphere in $\mathbb{R}^{3}$ is $4\pi h^{2}$). Thus, there exist constants $\mathfrak{C}_{1},\mathfrak{C}_{2}>0$ and an integer $h_{0}$ such that for all $h\geq h_{0}$,
$\mathfrak{C}_{1}\,h^{D-1}\leq|\mathcal{N}_{h}(v)|\leq\mathfrak{C}_{2}\,h^{D-1}.$ (53)
Proof of Theorem 6.2.
Assume that within $\mathcal{N}_{h}(v)$ there is a unique node $u^{*}$ with a strong influence $I(v,u^{*})=I^{*}>0$ and that for all other nodes $u\in\mathcal{N}_{h}(v)\setminus\{u^{*}\}$, the influence $I(v,u)$ is negligible ($\Delta\approx 0$). Then:
The total (or sum) influence is
$I_{\text{sum}}(v,h)=\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)\geq I(v,u^{*})=I^{*}.$ (54)
The mean influence is given by
$I_{\text{mean}}(v,h)=\frac{1}{|\mathcal{N}_{h}(v)|}\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)=\frac{1}{|\mathcal{N}_{h}(v)|}(I^{*}+\sum_{u\in\mathcal{N}_{h}(v)|u^{*}}I(v,u))=\frac{I^{*}+\Delta}{|\mathcal{N}_{h}(v)|}.$ (55)
Using the lower bound on $|\mathcal{N}_{h}(v)|$,
$I_{\text{mean}}(v,h)\leq\frac{I^{*}+\Delta}{\mathfrak{C}_{1}\,h^{D-1}}.$ (56)
Since $h^{D-1}\to\infty$ as $h\to\infty$ for $D\geq 2$, it follows that
$I_{\text{mean}}(v,h)\to 0,$ (57)
while $I_{\text{sum}}(v,h)\geq I^{*}$ remains non-vanishing. ∎
Proof of Corollary 6.3.
For a planar graph that is grid-like (for example, a two-dimensional lattice), set $D=2$. Then, $|\mathcal{N}_{h}(v)|=\Theta(h^{2-1})=\Theta(h).$ Repeating the same argument as above: the total influence satisfies $I_{\text{sum}}(v,h)\geq I^{*}.$ The mean influence is bounded by $I_{\text{mean}}(v,h)\leq\frac{I^{*}}{|\mathcal{N}_{h}(v)|}\leq\frac{I^{*}}{\mathfrak{C}_{1}\,h}.$ Hence, as $h\to\infty$, $I_{\text{mean}}(v,h)\to 0,$ which demonstrates that for a planar graph the dilution of the mean aggregated influence occurs at a rate proportional to $1/h$. ∎
Proof of Corollary 6.4.
Since $\mathcal{G}$ is grid-like in $D$ dimensions, the size of the aggregated $h$-hop neighborhood (the ball) grows as $|B_{h}(v)|=\Theta(h^{D}).$ Assuming that only one node $u^{*}$ in $B_{h}(v)$ has a significant influence $I^{*}$ while the influence of all other nodes is negligible, the total (or sum) influence satisfies $I_{\text{sum}}^{B}(v,h)\geq I^{*}$. Thus, the mean influence over $B_{h}(v)$ is $I_{\text{mean}}^{B}(v,h)\leq\frac{I^{*}}{\Theta(h^{D})}.$ That is, there exists a constant $\mathfrak{C}^{\prime}>0$ such that $I_{\text{mean}}^{B}(v,h)\leq\frac{I^{*}}{\mathfrak{C}^{\prime}\,h^{D}}.$ Since $h^{D}\to\infty$ as $h\to\infty$ for $D\geq 1$, it follows that $I_{\text{mean}}^{B}(v,h)\to 0.$
Proof of Corollary 6.5.
For any node $v\in V$, by the distinguished node assumption (reused from Theorem 6.2) there is at least one $u^{*}\in\mathcal{N}_{h}(v)$ with
$I(v,u^{*})\geq I^{*}.$ (58)
Since the total influence is a sum over nonnegative contributions, it follows that
$T_{h}(v)=\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)\geq I(v,u^{*})\geq I^{*}.$ (59)
Averaging over all nodes in $V$,
$\overline{T}_{h}=\frac{1}{|V|}\sum_{v\in V}T_{h}(v)\geq\frac{1}{|V|}\sum_{v\in V}I^{*}=I^{*}.$ (60)
This bound is independent of the size $|\mathcal{N}_{h}(v)|$ of the $h$-hop neighborhood and hence remains valid even as $|\mathcal{N}_{h}(v)|$ (or the overall number of nodes) tends to infinity.
$\overline{T}_{h}=\frac{1}{|V|}\sum_{v\in V}T_{h}(v)\geq I^{*},\quad\forall h.$ (61)